OpenStudy (anonymous):

Choose the correct simplification of (4a^4b^2)^2. Answer 16a^6b^4 8a^6b^4 16a^8b^4 8a^8b^4

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

I'm confused on what numbers i multiply

OpenStudy (accessdenied):

Here, we use these properties \( \large (mn)^x = m^x n^x\): \(\large (x^a)^b = x^{a\times b} \) First we bring the power outside to each factor, and then the second property for our variables.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

then what?

OpenStudy (accessdenied):

Simplify, and that should be it! The first step looks like this: \( \large (4a^4 b^2)^2 = 4^2 (a^4)^2 (b^2)^2 \) Kind of like an 'exponent-distributive property'.

OpenStudy (accessdenied):

The second property then lets us take those 'powers to a power' into one exponent: = 16 a^(4*2) b^(2*2) From here, we just simplify the multiplication: 4*2 = 8, 2*2 = 4...

OpenStudy (anonymous):

so last option?

OpenStudy (anonymous):

hey why does it say theres someone else viewing this but wont let me see their profile o_O

OpenStudy (anonymous):

its kind of like a facebook picture b4 u put ur own up...

OpenStudy (accessdenied):

4^2 = 4*4 = 16, so it can't be the last one with an 8 coefficient. It's a guest who hasn't registered. Hopefully they consider it. :P

OpenStudy (anonymous):

ahh ok

OpenStudy (anonymous):

first option?

OpenStudy (accessdenied):

Close... except I think you're confusing the multiplication of exponents with addition of exponents.... = 16 a^([[4*2]]) b^([[2*2]]) 4*2 = 8, 2*2 = 4; we can directly insert that back in: =16 a^8 b^4

OpenStudy (anonymous):

alright thanks